OpenStudy (anonymous):

(x + 4)(x − 4) Just wanna make shure this is right x^2-16?

hero (hero):

Yes, because (x + 4)(x - 4) = x(x - 4) + 4(x - 4) = x^2 - 4x + 4x + 16 = x^2 - 16

hero (hero):

It's called difference of squares

OpenStudy (anonymous):

ok sweet and then would (x + 4)(x + 4) be x^2+16

hero (hero):

Negative

OpenStudy (uri):

(x + 4)(x − 4) x(x-4)+4(x-4) x2-4x+4x-16 x2-16.

hero (hero):

(x + 4)(x + 4) = x(x + 4) + 4(x + 4) = x^2 + 4x + 4x + 16 = x^2 + 8x + 16

hero (hero):

It's called a perfect square trinomial.

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

i have to compare and contrast

OpenStudy (skullpatrol):

Any time you have the sum of two numbers times their difference, (a + b)(a - b) you get the difference of two squares: $$a^2-b^2$$

OpenStudy (anonymous):

so how to i compare and contrast x^2 - 16 and x^2 + 8x + 16

OpenStudy (skullpatrol):

The first is the difference of two squares: $$x^2 - 16=x^2-4^2$$

OpenStudy (anonymous):

ok

hero (hero):

One is a difference of squares, the other is a binomial square.

OpenStudy (skullpatrol):

The second is the square of a sum $$(a+b)^2$$

OpenStudy (skullpatrol):

$$(x + 4)^2$$

OpenStudy (anonymous):

so the similarity is the 8x

OpenStudy (skullpatrol):

There is no middle term in the difference of two squares.

OpenStudy (anonymous):

wait x^2 is a similarity

OpenStudy (skullpatrol):